2) A student performs this experiment and obtains a value of 3.02 x 10s s for the Rydberg Constant. Calculate the students
Answer #1

Actually ,

Rydberg's constant = 109678 cm​​​​​-1

The given frequency = 3.02 x 1015 s​​​​​​-1

frequency = Velocity of light / wavelength

Wavelength = Velocity of light / frequency

= 3 x 1010 cm s​​​​​​-1 / 3.02 x 1015 s​​​​​​-1

= 0.993 x 10-5 cm

1/ wavelength = 1/ 0.993 x 10-5 cm

= 1.007 x 105 cm​​​​​-1

= 100700 cm​​​​​-1​​​cm-1

This indicates Rydberg's constant

Now the difference ( error ) =( 109678 - 100700 ) cm​​​​​-1

= 8978 cm​​​​​-1

% error = ( Error / Theoretical value ) x 100

= (8978/ 109678 ) x 100

= 8.18

=

​​

=

​​​​